OpenStudy (anonymous):

For the polynomial shown below, find f(1). f(x)=3x^4-x^2+4x-2

OpenStudy (anonymous):

f(1) means to find the value of the function when x = 1. You have the name of the function, so what's its value when x = 1?

OpenStudy (ranga):

Put x = 1 f(1) = (3)(1^4) - (1^2) + (1) - 2 = (3)(1) - 1 + 1 - 2 = ?